# Overlay2025: 7th International Workshop on Artificial Intelligence and fOrmal VERification, Logic, Automata, and sYnthesis

### v2

@@ Abstract
Software correctness is ensured mathematically through formal verification, which involves the resources of generating formal requirement specifications and having an implementation that must be verified. Tools such as model-checkers and theorem provers ensure software correctness by verifying the implementation against the specification. Formal methods deployment is regularly enforced in the development of safety-critical systems e.g. aerospace, medical devices and autonomous systems. Generating these specifications from informal and ambiguous natural language requirements remains the key challenge. Our project, VERIFAI 111https://github.com/arshadbeg/OVERLAY2025_SupportingDocs.git, aims to investigate automated and semi-automated approaches to bridge this gap, using techniques from Natural Language Processing (NLP), ontology-based domain modelling, artefact reuse, and large language models (LLMs). This position paper presents a preliminary synthesis of relevant literature to identify recurring challenges and prospective research directions in the generation of verifiable specifications from informal requirements.
@@
The VERIFAI project [2] aims to address a central challenge of formal software engineering: translating informal, natural-language requirements into formal, verifiable specifications. This paper outlines the challenges and prospects in leveraging Large Language Models (LLMs) for formalising software requirements. The project aims to integrate techniques from Natural Language Processing (NLP), ontology-driven domain modelling, artefact reuse, and large language models (LLMs) to support the automated generation and traceability of formal specifications. Like all other fields, the development of large language models (LLMs) has opened a world of opportunities for the challenge of formalisation of software requirements. We consolidate initial findings to highlight research gaps and recurring difficulties in LLM-assisted formal specification generation.
@@
VERIFAI Research Framework: We propose our research framework, which integrates LLMs with NLP, ontology-driven modelling, and artefact reuse to support the semi-automated generation of verifiable formal specifications.
State-of-the-Art Synthesis: Through a focused literature review, we categorise and compare recent LLM-based tools and techniques—such as Req2Spec, SpecGen, AssertLLM, and nl2spec—highlighting their approaches, strengths, and limitations in requirement formalisation.
@@
Highlighting Gaps and Future Directions: Based on our synthesis, we outline critical open problems such as prompt instability, fragility of formal outputs, and the need for domain-specific context grounding. These pave the way for developing more robust LLM-based formal specification pipelines.
@@
Noting firstly that the main body of the paper is supported by several detailed Appendices. Section 2 describes a focused state of the art. Section 3 outlines the challenges and future directions for the research, supported by Appendix A that outlines the experiment setup and analysis performed on an example of the PathCrawler tool of Frama-C where we have re-simulated the methodology presented in [3]. This analysis is based on four provers available in Frama-C i.e. Alt-Ergo, Z3, CVC4 and CVC5. Appendix B presents the execution time comparison for these provers on the programs provided in [4]. Section 4 concludes the paper.
@@ 2 State of the Art
This section provides a focus on what we found to be the state of the art and its synthesis. The main research questions for conducting our systematic literature review on the topic were as follows:
RQ1: What methodologies leverage Large Language Models (LLMs) to transform natural language software requirements into formal notaions?
@@
The nl2spec tool supports interactive synthesis from unstructured requirements [8], while the SpecSyn tool improves sequence-to-sequence contract generation with a 21% accuracy gain [9]. Req2Spec converts 71% of BOSCH automotive requirements into formal specs [10], while SpecGen uses prompt mutation and verification feedback to improve LLM-generated specifications, succeeding on 279/384 benchmark programs [11]. In the hardware domain, AssertLLM synthesizes assertions with 89% correctness via multi-phase prompting and validation [12], while VLSI applications leverage LLMs for spec review and generation in SpecLLM [13]. Smart grid requirements have been formalised using GPT-4o and Claude 3.5, achieving F1 scores between 79% and 94% [14] The trend in F1-score observed by authors of [14] suggested that GPT-4o and Claude 3.5 not only maintain robustness but may actually benefit from increased system specification complexity, highlighting a potential alignment between model reasoning depth and specification richness. This aspect not mirrored in Gemini 1.5 or GPT-3.5-turbo and warranting further investigation. NASA’s software verification effort surfaced requirement errors, demonstrating practical utility of LLM-in-the-loop workflows [15].
NL-to-LTL translation has seen progress via few-shot prompting and dynamic reasoning [16], achieving 94.4% accuracy. Likewise, NL-to-JML contract synthesis for Java programs has been explored with promising results [17]. Historical systems like RSL [18], ARSENAL [19], and RML [20] demonstrated early rule-based and logic-based extraction pipelines, while hybrid neuro-symbolic systems offer greater reliability. SAT-LLM couples SMT solvers with LLMs to detect inconsistencies with F1 of 0.91 [21] and LeanDojo, ReProver, and Thor enhance formal proving via retrieval-augmented generation and LLM-guided reasoning [22, 23]. IDE-integrated efforts like those combining Copilot with PathCrawler and EVA demonstrate semi-automated ACSL spec generation [24, 3].
@@
Prompting strategies critically affect performance. While zero-shot prompting is strong in base performance [26], Chain-of-Thought (CoT) prompting improves logical flow via intermediate steps [27], though it suffers from context decay in long prompts (“lost-in-the-middle”) [28]. One-shot [29], few-shot [30], and SCoT [31] offer alternative trade-offs, yet zero-shot often remains competitive [32].
Advanced prompting methods like Automate-CoT generate CoT examples automatically [33], while Reprompting uses Gibbs sampling to escape prompt local optima [34]. Structured prompting with graphs and trees improves reasoning robustness and efficiency [35]. RAG (Retrieval-Augmented Generation) improves grounding for knowledge-intensive synthesis [36], while fine-tuning via LoRA enables lightweight domain adaptation [37].
@@
The effort accuracy is influenced by tool design and integration. For example, tools like nl2spec [8] improve generated specification quality through step-by-step refinement, adopting iterative and user-in-the-loop approach to help address some LLM limitations. Similarly, prompt engineering techniques utilising guided templates or Chain-of-Thought (CoT) [7, 40, 41] promised improved output coherence and correctness. But again, these strategies work well in scenarios involving localised tasks, such as assertion synthesis or narrow-scope descriptions. As the program size and complexity of the specification goal increase, the chances of ambiguity, under-specification, and logical inconsistency increase. Therefore, we can conclude that current LLM architectures excel in focused, declarative tasks but require augmentation for broader specification goals.
We can conclude from our synthesis of the current literature that the research trend is increasing in combining the strength of LLMs, symbolic reasoning and iterative user interaction. At the moment, assertion generation dominates in terms of accuracy and usability, but the parallel efforts of better prompt design, domain-specific fine-tuning, and verifier-in-the-loop are closely matching the performance of the process. The challenges of abstraction and consistency drive research efforts in this domain.
@@ 3 Challenges and Future Directions
Semantic ambiguity (C1) due to natural language remains a critical issue, needing structured domain knowledge and improved human-in-the-loop interventions. The lack of publicly available, high-quality datasets (C2) hinders model training, reproducibility, and scalability. Tool interoperability (C3) is impeded by incompatible formats and absence of standardised interfaces, complicating automation. Ensuring traceability across artefact lifecycles (C4) is essential but difficult without explainable and collaborative processes. Lastly, explainability and user trust (C5) are limited by opaque model behavior and insufficient rationale in outputs. To address these, human-in-the-loop formalisation (F1) offers controlled semi-automation and improved trust, while multi-modal artefact alignment (F2) enables contextual completeness via diverse input formats. The creation of standardised benchmarks (F3) would mitigate dataset-related limitations and promote progress. Neuro-symbolic reasoning (F4) blends LLM flexibility with logic-based precision, enhancing model reliability. Finally, interactive traceability tools (F5) that support collaboration, visual navigation, and auditability are crucial for regulated and complex software domains.
@@ Appendix A Work in progress and initial experiment setup
As initial experiment setup, we re-simulated the methodology presented in [3] by applying it to TriType.c program, which was selected from the online interface of the PathCrawler tool available in the Frama-C ecosystem. The program was analysed using the original workflow – combining Large Language Models (LLMs) with symbolic outputs from PathCrawler and EVA – to generate ACSL specifications. The use of PathCrawler provided concrete path-based input/output examples that guided the LLM toward generating more context-relevant and semantically aligned annotations. Figure 1 shows the output of the Pathcrawler tool. Our plan is to conduct experiments with simple and complex programs, representing a diverse set of procedural constructs and control flows, allowing us to evaluate the methodology across a realistic and varied set of inputs. We have created a public repository on GitHub https://github.com/arshadbeg/OVERLAY2025_SupportingDocs.git to add resources.

### v3

@@ Abstract
Software correctness is ensured mathematically through formal verification, which requires the generation of a formal requirement specification and an implementation that must be verified. Tools such as model-checkers and theorem provers ensure software correctness by verifying the implementation against the specification. Formal methods deployment is regularly enforced in the development of safety-critical systems e.g. aerospace, medical devices and autonomous systems. Generating these specifications from informal and ambiguous natural language requirements remains the key challenge. Our project, VERIFYAI 111https://github.com/arshadbeg/OVERLAY2025_SupportingDocs.git, aims to investigate automated and semi-automated approaches to bridge this gap, using techniques from Natural Language Processing (NLP), ontology-based domain modelling, artefact reuse, and large language models (LLMs). This position paper presents a preliminary synthesis of relevant literature to identify recurring challenges and prospective research directions in the generation of verifiable specifications from informal requirements.
@@
The VERIFYAI project [2] aims to address a central challenge of formal software engineering: translating informal, natural-language requirements into formal, verifiable specifications. This paper outlines the challenges and prospects in leveraging Large Language Models (LLMs) for formalising software requirements. The project aims to integrate techniques from Natural Language Processing (NLP), ontology-driven domain modelling, artefact reuse, and large language models (LLMs) to support the automated generation and traceability of formal specifications. Like all other fields, the development of large language models (LLMs) has opened a world of opportunities for the challenge of formalisation of software requirements. We consolidate initial findings to highlight research gaps and recurring difficulties in LLM-assisted formal specification generation. Noting firstly that the main body of the paper is supported by several detailed Appendices available at https://github.com/arshadbeg/OVERLAY2025_SupportingDocs.git.
@@
VERIFYAI Research Framework: We propose our research framework, which integrates LLMs with NLP, ontology-driven modelling, and artefact reuse to support the semi-automated generation of verifiable formal specifications.
State-of-the-Art Synthesis: Through a focused literature review, we categorise and compare (Section 2, supported by Appendix A) recent LLM-based tools and techniques—such as Req2Spec, SpecGen, AssertLLM, and nl2spec—highlighting their approaches, strengths, and limitations in requirement formalisation.
@@
Highlighting Gaps and Future Directions: Based on our synthesis, we outline critical open problems such as prompt instability, fragility of formal outputs, and the need for domain-specific context grounding (Appendix D). These pave the way for developing more robust LLM-based formal specification pipelines.
@@
The structure of the paper is: Section 2 describes a focused state of the art. Section 3 outlines the challenges and future directions for the research, supported by Appendix A that outlines the experiment setup and analysis performed on an example of the PathCrawler tool of Frama-C where we have re-simulated the methodology presented in [3]. This analysis is based on four provers available in Frama-C i.e. Alt-Ergo, Z3, CVC4 and CVC5. Appendix B presents the execution time comparison for these provers on the programs provided in [4]. Section 4 discusses our plans for future work and Section 5 concludes the paper.
@@ 2 State of the Art
This section synthesises what we found to be the state of the art. The main research questions for conducting our systematic literature review on the topic were as follows:
RQ1: What methodologies leverage Large Language Models (LLMs) to transform natural language software requirements into formal notations?
@@
Here, we summarise our main findings. For a comprehensive overview of the literature survey, including detailed comparisons and categorised insights, we encourage readers to consult the accompanying GitHub repository, which presents the full set of supporting tables.
@@
The nl2spec tool supports interactive synthesis from unstructured requirements [8], while the SpecSyn tool improves sequence-to-sequence contract generation with a 21% accuracy gain [9]. Req2Spec converts 71% of BOSCH automotive requirements into formal specs [10], while SpecGen uses prompt mutation and verification feedback to improve LLM-generated specifications, succeeding on 279 out of 384 benchmark programs [11]. In the hardware domain, AssertLLM synthesizes assertions with 89% correctness via multi-phase prompting and validation [12], while VLSI applications leverage LLMs for spec review and generation in SpecLLM [13]. Smart grid requirements have been formalised using GPT-4o and Claude 3.5, achieving F1 scores between 79% and 94% [14]. The trend in F1-scores observed by authors of [14] suggested that GPT-4o and Claude 3.5 not only maintain robustness but may actually benefit from increased system specification complexity, highlighting a potential alignment between model reasoning depth and specification richness. This aspect not mirrored in Gemini 1.5 or GPT-3.5-turbo and warranting further investigation. NASA’s software verification effort surfaced requirement errors, demonstrating practical utility of LLM-in-the-loop workflows [15].
NL-to-LTL translation has seen progress via few-shot prompting and dynamic reasoning [16], achieving 94.4% accuracy. Likewise, NL-to-JML contract synthesis for Java programs has been explored with promising results [17]. Historical systems like RSL [18], ARSENAL [19], and RML [20] demonstrated early rule-based and logic-based extraction pipelines, while hybrid neuro-symbolic systems offer greater reliability. SAT-LLM couples SMT solvers with LLMs to detect inconsistencies with F1 of 0.91 [21] and LeanDojo, ReProver, and Thor enhance formal proving via retrieval-augmented generation and LLM-guided reasoning [22, 23]. IDE-integrated efforts like those combining Copilot with PathCrawler and EVA demonstrate semi-automated ACSL specification generation [24, 3].
@@
LLM selection and prompting strategies critically affect performance. While zero-shot prompting is strong in base performance [26], one-shot [27] and few-shot [28] offer alternative trade-offs. Chain-of-Thought (CoT) prompting improves logical flow via intermediate steps [29], and like Structured CoT (SCoT) [30] it can suffer from context decay in long prompts (“lost-in-the-middle”) [31], yet zero-shot often remains competitive [32].
Advanced prompting methods like Automate-CoT generate CoT examples automatically [33], while Reprompting uses Gibbs sampling to escape prompt local optima [34]. Structured prompting with graphs and trees improves reasoning robustness and efficiency [35]. RAG (Retrieval-Augmented Generation) improves grounding for knowledge-intensive synthesis [36]. [37] discusses a wider range of almost 30 prompting strategies, some of which seem not to have been explored in relation to formalising specifications. Of course, this does not include related approaches such as fine-tuning LLM via LoRA adaptation training [38], but this may only be applicable when there is access to the LLM’s architecture and weights. Additionally, reinforcement learning (RL) may help with specific challenges, opening even more avenues for exploration.
@@
The effort accuracy is influenced by tool design and integration. For example, tools like nl2spec [8] improve generated specification quality through step-by-step refinement, adopting iterative and user-in-the-loop approach to help address some LLM limitations. Similarly, prompt engineering techniques utilising guided templates or Chain-of-Thought (CoT) [7, 41, 42] promised improved output coherence and correctness. These strategies work well in scenarios involving localised tasks, such as assertion synthesis or narrow-scope descriptions. As the program size and complexity of the specification goal increase, the chances of ambiguity, under-specification, and logical inconsistency increase. Therefore, we conclude that the current LLM architectures excel in focused, declarative tasks but require augmentation for broader specification goals. However, [43] showed that different versions of language models including LLM, can vary greatly in their responses to the same queries, suggesting that much experimental work might be required to achieve optimum results.
We conclude from our synthesis of the current literature that the research trend is increasing in combining the strength of LLMs, symbolic reasoning and iterative user interaction. At the moment, assertion generation dominates in terms of accuracy and usability, but the parallel efforts of better prompt design, domain-specific fine-tuning, and verifier-in-the-loop are closely matching the performance of the process. The challenges of abstraction and consistency drive research efforts in this domain.
@@ 3 Challenges and Future Directions
Semantic ambiguity (C1) due to natural language remains a critical issue, needing structured domain knowledge and improved human-in-the-loop interventions. The lack of publicly available, high-quality datasets (C2) hinders model training, reproducibility, and scalability. Tool interoperability (C3) is impeded by incompatible formats and absence of standardised interfaces, complicating automation. Ensuring traceability across artefact lifecycles (C4) is essential but difficult without explainable and collaborative processes. In addition, explainability and user trust (C5) are limited by opaque model behavior and insufficient rationale in outputs. To address these, human-in-the-loop formalisation (F1) offers controlled semi-automation and improved trust, while multi-modal artefact alignment (F2) enables contextual completeness via diverse input formats. The creation of standardised benchmarks (F3) would mitigate dataset-related limitations and promote progress. Neuro-symbolic reasoning (F4) blends LLM flexibility with logic-based precision, enhancing model reliability. Finally, interactive traceability tools (F5) that support collaboration, visual navigation, and auditability are crucial for regulated and complex software domains.
4 Embedding LLMs in the Specification Generation
4.1 Planned Evaluation of Prompting Strategies
A systematic and quantitative evaluation of prompting strategies is the central part of our planned research. In particular, we intend to compare zero-shot, one-shot, few-shot, and Chain-of-Thought (CoT) prompting across both assertion-level and full-contract generation tasks. Prior work in the literature [26–35] suggests that prompt type and articulation can significantly influence specification quality. However, a comprehensive evaluation requires a larger and more diverse benchmark than we currently report.
Future experiments will therefore assess prompt sensitivity using precision, recall, and F1-based correctness metrics, and investigate robustness under small variations of prompt formulation. We anticipate that assertion-level tasks will prove more stable under prompt rephrasing, whereas full-contract synthesis may show higher variability — an observation that motivates deeper analysis in follow-up work. Our current contribution is to highlight the importance of prompt design in our work and to outline how this dimension will be systematically investigated going forward.
4.2 Human-in-the-Loop Integration and Interoperability Considerations
Our experiments currently implement the early stages of a human-in-the-loop process through manual review. In both the Tritype baseline and PathCrawler-augmented workflows, every LLM-generated ACSL specification was checked by at least one author with formal methods expertise. This review ensured (i) semantic alignment with intended behavior, (ii) logical completeness, and (iii) iterative refinement by feeding corrected fragments back into the prompts. Although active learning is not yet integrated, our project approach anticipates it: revised specifications, along with their source code and verification results, can be versioned and selectively reused for retraining or fine-tuning.
Figure 1(b) shows our proposed workflow based on the state of the art and figured out challenges in Sections 2 and 3. Natural language requirements and domain ontologies form the input, grounding the LLM in the target context. Different prompt strategies (zero-shot, few-shot, and Chain-of-Thought) shape how inputs are presented for specification generation. Outputs are stored in a tool-neutral intermediate format (JSON-LD), which can be translated into the syntax required by verification tools. These tools check the generated specifications, while symbolic reasoning provides constraint-based feedback that helps refine them. Human reviewers remain part of the loop, validating results, focusing on manageable units, and collecting useful examples for future adaptation. The design is modular, so new reasoning engines, prompt methods, or domain adapters can be added without changing the overall pipeline.
For interoperability, the pipeline currently uses custom scripts to convert LLM outputs to tool-specific formats (e.g., ACSL for Frama-C, JML for OpenJML). While functional, this approach is not generalisable. As part of VERIFYAI, we plan to design a lightweight JSON-LD-based schema that maps to multiple formal languages. This would allow LLM outputs to be stored in a tool-agnostic format and exported to different targets, potentially supporting Frama-C, OpenJML, Dafny, and others with minimal per-tool changes.
4.3 Symbolic Reasoning, Traceability, and Scalability Outlook
In our initial Tritype experiments, symbolic solvers were used only for post-generation verification. We see potential for tighter integration, where solver feedback—such as unsatisfiable path conditions—could guide the LLM during generation, reducing logical errors before verification. Currently, traceability relies on manual annotations linking specification fragments to source code and natural-language requirements. We are exploring semi-automated approaches where the LLM proposes initial links for human validation. These links could be stored in a graph database to support version-aware navigation and visual trace maps, which would be particularly valuable in regulated settings.
We aim to build datasets that combine real-world, diverse requirements with verified specifications and execution traces. Currently, we are curating a small seed set from open-source safety-critical software, supplemented with synthetic examples to cover edge cases. While synthetic data is useful, it lacks the nuances of industrial requirements, so mixed datasets appear most promising. VERIFYAI’s prototype handles single-module programs well, but multi-module systems pose memory and latency challenges. To manage complexity, we plan to employ hierarchical specification synthesis, verifying each module independently prior to integration. Additionally, we aim to release an annotated subset of the dataset to support transparency and reproducibility.
@@ Appendix A Work in progress and initial experiment setup
As initial experiment setup, we re-simulated the methodology presented in [3] by applying it to TriType.c program, which was selected from the online interface of the PathCrawler tool available in the Frama-C ecosystem. The program was analysed using the original workflow – combining Large Language Models (LLMs) with symbolic outputs from PathCrawler and EVA – to generate ACSL specifications. The use of PathCrawler provided concrete path-based input/output examples that guided the LLM toward generating more context-relevant and semantically aligned annotations. Figure 2 shows the output of the Pathcrawler tool. Our plan is to conduct experiments with simple and complex programs, representing a diverse set of procedural constructs and control flows, allowing us to evaluate the methodology across a realistic and varied set of inputs. A repository of related resources is available at https://github.com/arshadbeg/OVERLAY2025_SupportingDocs.git.